Lana ordered concert tickets that cost $7.50 for children and $12.00 for adults. She ordered 8 more children's tickets than adults' tickets. Her total bill was $138. How many of each type of ticket did she order?
Having problems setting up and solving. Please help

Answer by sofiyac(983)

lets say she got x of adult tickets, then children's would be x+8
7.50(x+8)+12x=138
7.50x+60+12x=138
19.50x=78
x=4 so we have 4 adult tickets and 4+8=12 children's tickets